Sharon has some one-dollar bills and some five dollar bills. She has 14 bills. The value of the bills is $30. Solve a system of equations using elimination to find how many of each kind of bill she has.
Answer by mananth(12270)

let 1 dollar bills be x
let 5 dollar bills be y
x+y=14-------------1
x+5y=30------------2
subtract
x-y - ( x+5y)= 14-30
x+y-x-5y=-16
-4y=-16
y=4 ----- 5 dollar bills
She has 10 1 dollar bills